# Is Compression Really Linear with Code Intelligence?

### v4

@@ 1 Introduction
Data compression seeks to represent information using fewer bits than its original encoding by exploiting statistical redundancies bookstein1990compression ; johnson2003introduction ; rissanen2000coding ; yvinec2021red . According to Kolmogorov complexity theory ruffini2007information ; li2008introduction ; yoran2025the , effective compression relies upon identifying latent regularities and structural features within data, a process fundamental to intelligence and its core cognitive abilities such as learning, comprehension, and generalization zanto2011causal ; bargmann2014brain . Recently, the next-token prediction objective, central to the pre-training of LLMs, has been recognized as mirroring compression principles deletang2024language . This has led to viewing LLMs as powerful compressors and exploring compression-based metrics for evaluating their learned capabilities deletang2024language ; wang2024language ; huang2025look ; huang2024compression . However, code as a unique data modality differs from natural language in its formal syntactic structures and abstraction mechanisms 294586 ; zhu-etal-2024-survey ; parțachi2024bringing . For instance, information missing in code can have far more severe consequences aryabumi2024code . Therefore, a dedicated investigation into code compression and code intelligence is of paramount importance gnvv2023the .
@@
Figure 1: (a) shows the linear relationship between code BPC and code intelligence as claimed in huang2024compression . (b) illustrates the linear fitting result of advanced Code LLMs on the same setting.
@@
The previous study huang2024compression proposes the linear relationship between code compression and code intelligence, using bits per character (BPC) as a proxy for the former and average performance on three Python code completion benchmarks: HumanEval chen2021evaluating , MBPP austin2021program , and DS-1000 lai2023ds . Furthermore, evaluated LLMs are earlier versions, which possessed relatively weaker code capabilities and had seen limited adoption in current applications. Our preliminary evaluation result illustrated in Figure 1(b) reveals no clear correlation between the average performance of several modern code LLMs and their code BPC. Code intelligence spans numerous programming languages (e.g., Python, C++, and Java) and encompasses several fundamental “atomic” tasks - code understanding (explain), reasoning, generation, and repair chen2021evaluating . Consequently, evaluating code intelligence solely through Python code completion is insufficiently comprehensive. Furthermore, code intelligence has seen significant progress with the emergence of advanced Code LLMs. These LLMs are pre-trained on substantially larger code corpora, exhibit superior performance across a variety of coding tasks, and are increasingly utilized in real-world development environments. This necessitates a more in-depth study of the relationship between code compression and code intelligence.
@@
Even with a well-defined set of tasks and models, evaluating the code intelligence of pretrained LLMs remains a challenge. They often struggle with tasks beyond simple code completion. For instance, they may generate code with non-standard formatting or misinterpret task requirements raychev2014code . Common evaluation paradigms like few-shot prompting reynolds2021prompt ; ye2022unreliability heavily rely on the choice and quality of demonstrations, exhibit limited generalization, and require significant effort in designing specific prompts for each distinct task. Furthermore, some LLMs incorporate instruction-tuning data during their pre-training phase, which artificially inflates performance on benchmarks. Illustratively, Figure 1(b) highlights inconsistencies in the self-reported performance of various Code LLMs. These factors highlight an urgent need for more robust and fair evaluation methods capable of revealing the genuine capabilities of pre-trained Code LLMs.
@@
To address these pressing issues, we undertake a two-pronged approach. Firstly, to establish a comprehensive scope for code intelligence, we have curated a diverse suite of open-source Coder LLMs, encompassing prominent families such as Qwen Coder hui2024qwen2 , Code Llama roziere2023code , Starcoder li2023starcoder and Opencoder huang2024opencoder . Furthermore, we propose OmniCodeBench, which spans multiple programming languages and diverse code-related tasks. Secondly, to ensure a fair and veridical assessment of pre-trained LLMs, we introduce Format Annealing, which is a lightweight, controlled training method on a consistent and publicly available dataset. This process is designed to elicit the intrinsic ability of each model in a standardized manner, reflecting their true intelligence. Subsequently, to quantify compression efficacy, we have meticulously constructed a large-scale, high-quality code validation set derived from GitHub, ensuring it is previously unseen by the models, diverse in its content, and unbiased in its composition. Using this validation set, we measure the bits-per-character (BPC) of each LLM. The empirical results reveal a significant departure from previous findings: Code Intelligence exhibits a logarithmic relationship with Code BPC. We posit that earlier observations of linearity likely stemmed from analyses focused on higher compression regimes, where the tail of a logarithmic curve can indeed approximate linearity.
Our contributions are summarized as below:
@@
  • We establish a robust and comprehensive framework for investigating the relationship between code intelligence and compression by selecting high-quality, multi-task, and multi-language code benchmarks OmniCodeBench, curating a diverse set of contemporary open-source Coder LLMs, and constructing a novel, large-scale code validation set.
  • We introduce Format Annealing, a principled and transparent methodology for eliciting and fairly evaluating the performance of pre-trained models on these benchmarks using publicly accessible data.
  • Our empirical investigation reveals a fundamental logarithmic relationship between code compression rate (BPC) and Code Intelligence, offering a new perspective that refines and contextualizes previous findings of linearity.
@@ 2.1 Compression as Metric
Early evaluations of LLM intelligence primarily relied on averaged performance across various benchmarks brown2020language . However, the scope of benchmarks was constrained by the pretrained LLMs’ nascent instruction-following capabilities. Recognizing the intrinsic consistency between pre-training objectives and information compression, researchers began to investigate the relationship between model compression rates and intelligence levels. wei2023skywork first reveals the linear correlation between compression ratios and downstream benchmark metrics. Subsequent studies huang2024compression further substantiated the persistence of this linearity across diverse model series and benchmark domains. Despite these advancements, prior work has predominantly treated code as a subdomain of natural language, which overlooks fundamental multi-dimensional distinctions. Code adheres to strict, formal syntactic rules, precluding the ambiguities and flexible grammatical structures inherent in natural language zhu-etal-2024-survey . Furthermore, abstraction in code manifests through explicit, formally-defined constructs such as classes, functions, and APIs, contrasting sharply with the predominantly cognitive and conceptual nature of abstraction in natural language de2003side ; abbott2008abstraction . These critical differences motivate a dedicated inquiry into whether the observed relationship between compression and intelligence in natural language extends uniformly to code.
@@ 2.2 Code Intelligence
The remarkable capabilities of LLMs in code-related tasks have spurred the concept of “code intelligence”—the comprehensive ability to understand, generate, and manipulate code (rabin2021understanding, ; wan2024deep, ). Early evaluations typically focused on Python code completion tasks, such as HumanEval and MBPP (zhu-etal-2024-survey, ). However, as code applications and workflows grow in complexity (wang2024agents, ; roberttransforming, ), it is increasingly evident that a robust notion of code intelligence must extend beyond unilingual assessments to encompass proficiency across multiple programming languages (cassano2022multipl, ; peng-etal-2024-humaneval, ; chai2024mcevalmassivelymultilingualcode, ; liu2024mdeval, ). Furthermore, a growing body of research suggests that comprehensive code intelligence is not a monolithic entity but consists of a set of fundamental, or “atomic”, tasks. While existing taxonomies for code-related LLM tasks vary—for instance, by input/output modalities (e.g., NL-PL, PL-NL) or by user intent (chen2024survey, ) —they converge on a core set of underlying capabilities. We identify these as: (1) Code Generation: automated creation of source code from natural language descriptions (chen2021evaluating, ; austin2021program, ; zhuo2024bigcodebench, ). (2) Code Explanation (or Understanding): grasp and articulate the syntax, semantics, functionality, and underlying intent of a given code snippet (meyerson2025position, ; fang2024large, ; nam2024using, ). (3) Code Reasoning: infer code properties and predict code behavior without execution (dehghan2024assessing, ; chen2024reasoning, ; li2025system, ). (4) Code Repair: identify and generate patches to correct bugs, vulnerabilities, or other errors in source code (pasquale2025challenges, ; liu2024mdeval, ). While “code intelligence” is a broad concept, its decomposition into these four atomic tasks provides a more structured and actionable framework for quantitative evaluation.
@@ 3 Code Compression as Code Intelligence
We investigate the relationship between compression ratio and code intelligence. In § 3.1, we propose sliding window BPC for quantifying compression ratio and construct a set of robust validation corpus. In § 3.2, we present OmniCodeBench, a multi-lingual and multi-task code benchmark for comprehensive evaluation of code intelligence capabilities of LLMs.
@@ 3.1.1 Compression Ratio Criteria
$BPC=\sum_{i=1}^{N}\log\left(\frac{1}{p_{i}}\right)\cdot\frac{1}{M}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left(-\log(p_{i})\right)}_{\mathrm{train~{}loss}}\cdot\frac{N}{M}=R_{1}\cdot R_{2}=R$ (1)
where $M$, $N$ denote corpus size and vocabulary size, respectively. $R_{1}$, the first term of the BPC formula, captures model compression through next-token prediction loss, which is equivalent to the training loss upon first exposure. $R_{2}$, the second term of the BPC formula, represents vocabulary compression rate. It is worth noting that due to differences in vocabulary size, cross-family model comparisons need to consider $R_{1}\cdot R_{2}$, while analysis within the same family can focus solely on $R_{1}$.
@@ 3.1.2 Sliding Window BPC
A critical yet frequently overlooked factor in vanilla BPC evaluation lies in the heterogeneity of models’ context window lengths $W$. Previous practices often simply truncate validation samples exceeding $W$, thereby discarding cross-chunk contextual dependencies and introducing bias when comparing models with divergent context length. To address this, we propose a sliding window protocol with stride adaptation for precise BPC calculation, as shown in Algorithm 1.
Algorithm 1 Sliding Window BPC
1:Initialize position $t\leftarrow 0$, stride $V\leftarrow W/4$
2:while $t+W\leq M$ do
3:     Feed $x_{t:t+W}$ to LLM
4:     if $t=0$ then $\triangleright$ First window
5:         Capture $\log p_{\theta}(x_{i}|x_{<i})$ for $\forall i\in[0,W)$
6:     else
7:         Capture $\log p_{\theta}(x_{i}|x_{<i})$ for $i\in[t+W-V,t+W)$
8:     end if
9:     $t\leftarrow t+V$
10:end while
@@
Figure 2: Programming language token distribution in the open-source and proprietary validation set. In (b) The unlabeled categories Perl, Shell, C++, Vue, and SQL account for 0.88%, 0.80%, 0.66%, 0.38%, and 0.03% respectively.
@@ 3.1.3 Robust Validation Corpus Construction
Following the aforementioned criteria, we construct two high-quality development corpora, one based on proprietary internal data and the other on publicly available open-source data.
@@
Public corpus   To mitigate limitations in linguistic diversity and distributional biases inherent in proprietary datasets, we additionally construct an open-source validation set. We select GitHub as our data source and employ timestamp-based filtering to prevent validation data leakage. Specifically, based on the cutoff date of existing code LLM pre-training data, we collect all GitHub repositories created between May 2024 and November 2024. Subsequently, we follow OpenCoder’s (huang2024opencoder, ) pipeline to perform filtration, removing non-code data and deduplicating the dataset. To minimize bias, we employ rule-based filtering rather than model-based filtering approaches. Additionally, we parse the code to confirm its compilability and apply heuristic-based selection (e.g., filtering by natural language comment length) to preserve high-quality data. However, due to potential historical code leakage from GitHub branches, we adopt MinHash-LSH algorithm to prevent the inclusion of duplicate or near-duplicate code in the validation set. To ensure balanced domain and language distribution in the development set, we perform weighted sampling by repository category, mitigating bias toward high-resource languages. The distribution of validation set is illustrated in Figure 2.
@@ 3.2.1 OmniCodeBench
The evaluation of code intelligence in LLMs remains an open research question. While existing studies predominantly rely on Python-specific benchmarks like HumanEval (chen2021evaluating, ) and MBPP (austin2021program, ), Python alone cannot represent the full spectrum of programming concepts luo2024python . Our evaluation, therefore, encompasses multiple programming languages. Moreover, conventional assessments focus primarily on the code completion task, which constitutes merely a fraction of real-world coding intelligence. We address this limitation through a comprehensive multi-task evaluation framework. To this end, we propose OmniCodeBench, a multi-lingual and multi-task code benchmark capable of holistic and robustly evaluating code intelligence. Concretely, OmniCodeBench consists of McEval chai2024mceval for multi-lingual code generation, BIRD li2023can and Spider yu2018spider for SQL code generation, LiveCodeBench jain2024livecodebench for program repair, HumanEvalExplain muennighoff2023octopack for code explanation, CRUXEval gu2024cruxeval for code reasoning, and HumanEval chen2021evaluating , MBPP austin2021program , BigCodeBench zhuo2024bigcodebench .
@@
StarCoder2-3B (lozhkov2024starcoder, ) 136 1
CodeGemma-2B (team2024codegemma, ) 55 2
CodeLlama-7B (roziere2023code, ) 10 2
Qwen2.5-Coder-7B (hui2024qwen2, ) 13 4
Granite-8B-Code (mishra2024granite, ) 12 3
@@ 3.2.2 Format Annealing
Code-related tasks require extracting code from the model’s responses and verifying its correctness through test cases. However, the response format of the model is often difficult to perfectly cover, leading to omissions during extraction. Additionally, the model frequently outputs meaningless placeholders, such as todo and pass, severely impacting the accurate evaluation of code intelligence. This phenomenon is caused by the excessive occurrence of related patterns in the pre-training data huang2024opencoder . To address this, we construct an annealed training set to ensure: (1) preventing external knowledge infusion beyond the original capacity of foundation models, and (2) maintaining strict format compliance with downstream benchmarks. Following these requirements, we integrate three data sources: CommitPackFT muennighoff2023octopack , CodeSearchNet husain2019codesearchnet , and LeetCode (huang2024opencoder, ). All three datasets originate from GitHub, which is the major component of the pretraining corpora of Code LLMs. We run the data cleaning pipelines of various Code LLMs on this data, ensuring the remaining parts were seen during pre-training. Subsequently, all non-functional placeholders are removed, and the content is reformatted according to the prompts of downstream benchmarks, ensuring alignment for evaluation. The final curated dataset contains 4,995 training samples. The complete data synthesis pipeline for each dataset is presented in Appendix.
The curated dataset is then used for the specialized continual training phase for all models. During this phase, after each epoch, we leverage the MBPP+ evalplus to assess the proportion of post-processed outputs that are empty strings. Training for each model is terminated when this proportion drops to 1%. This early stopping strategy is employed to minimize alterations to the models’ original parameter distributions. Table 1 shows the comparison of the number of empty strings before and after format annealing for five models. It shows that format annealing effectively mitigates the inherent limitations of foundational models in format compliance, ensuring fair comparison through format-controlled evaluation.

### v5

# Title for citation

@@ Abstract
keywords:
large language model, code compression, bit per character, benchmark, scaling law
@@ 1 Introduction
Data compression seeks to represent information using fewer bits than its original encoding by exploiting statistical redundancies [6, 29, 48, 60]. According to Kolmogorov complexity theory [51, 33, 58], effective compression relies upon identifying latent regularities and structural features within data, a process fundamental to intelligence and its core cognitive abilities such as learning, comprehension, and generalization [61, 5]. Recently, the next-token prediction objective, central to the pre-training of LLMs, has been recognized as mirroring compression principles [16]. This has led to viewing LLMs as powerful compressors and exploring compression-based metrics for evaluating their learned capabilities [16, 54, 24, 25]. However, code as a unique data modality differs from natural language in its formal syntactic structures and abstraction mechanisms [22, 62, parțachi2024bringing]. For instance, information missing in code can have far more severe consequences [3]. Therefore, a dedicated investigation into code compression and code intelligence is of paramount importance [18].
@@
Figure 1: (a) shows the linear relationship between code BPC and code intelligence as claimed in [25]. (b) illustrates the linear fitting result of advanced Code LLMs on the same setting.
@@
The previous study [25] proposes the linear relationship between code compression and code intelligence, using bits per character (BPC) as a proxy for the former and average performance on three Python code completion benchmarks: HumanEval [13], MBPP [4], and DS-1000 [30]. Furthermore, evaluated LLMs are earlier versions, which possessed relatively weaker code capabilities and had seen limited adoption in current applications. Our preliminary evaluation result illustrated in Figure 1(b) reveals no clear correlation between the average performance of several modern code LLMs and their code BPC. Code intelligence spans numerous programming languages (e.g., Python, C++, and Java) and encompasses several fundamental “atomic” tasks - code understanding (explain), reasoning, generation, and repair [13]. Consequently, evaluating code intelligence solely through Python code completion is insufficiently comprehensive. Furthermore, code intelligence has seen significant progress with the emergence of advanced Code LLMs. These LLMs are pre-trained on substantially larger code corpora, exhibit superior performance across a variety of coding tasks, and are increasingly utilized in real-world development environments. This necessitates a more in-depth study of the relationship between code compression and code intelligence.
@@
Even with a well-defined set of tasks and models, evaluating the code intelligence of pretrained LLMs remains a challenge. They often struggle with tasks beyond simple code completion. For instance, they may generate code with non-standard formatting or misinterpret task requirements [46]. Common evaluation paradigms like few-shot prompting [47, 57] heavily rely on the choice and quality of demonstrations, exhibit limited generalization, and require significant effort in designing specific prompts for each distinct task. Furthermore, some LLMs incorporate instruction-tuning data during their pre-training phase, which artificially inflates performance on benchmarks. Illustratively, Figure 1(b) highlights inconsistencies in the self-reported performance of various Code LLMs. These factors highlight an urgent need for more robust and fair evaluation methods capable of revealing the genuine capabilities of pre-trained Code LLMs.
@@
To address these pressing issues, we undertake a two-pronged approach. Firstly, to establish a comprehensive scope for code intelligence, we have curated a diverse suite of open-source Coder LLMs, encompassing prominent families such as Qwen Coder [26], Code Llama [50], Starcoder [34] and Opencoder [23]. Furthermore, we propose OmniCodeBench, which spans multiple programming languages and diverse code-related tasks. Secondly, to ensure a fair and veridical assessment of pre-trained LLMs, we introduce Format Annealing, which is a lightweight, controlled training method on a consistent and publicly available dataset. This process is designed to elicit the intrinsic ability of each model in a standardized manner, reflecting their true intelligence. Subsequently, to quantify compression efficacy, we have meticulously constructed a large-scale, high-quality code validation set derived from GitHub, ensuring it is previously unseen by the models, diverse in its content, and unbiased in its composition. Using this validation set, we measure the bits-per-character (BPC) of each LLM. The empirical results reveal a significant departure from previous findings: Code Intelligence exhibits a logarithmic relationship with Code BPC. We posit that earlier observations of linearity likely stemmed from analyses focused on higher compression regimes, where the tail of a logarithmic curve can indeed approximate linearity.
@@
Our contributions are summarized as below: 1) We establish a robust and comprehensive framework for investigating the relationship between code intelligence and compression by selecting high-quality, multi-task, and multi-language code benchmarks to construct OmniCodeBench, curating a diverse set of contemporary open-source Coder LLMs, and constructing a novel, large-scale code validation set. 2) We introduce Format Annealing, a principled and transparent methodology for eliciting and fairly evaluating the performance of pre-trained models on these benchmarks using publicly accessible data. 3) Our empirical investigation reveals a fundamental logarithmic relationship between code compression rate and Code Intelligence, offering a new perspective that refines and contextualizes previous findings of linearity.
@@ 2.1 Compression as Metric
Early evaluations of LLM intelligence primarily relied on averaged performance across various benchmarks [7]. However, the scope of benchmarks was constrained by the pretrained LLMs’ nascent instruction-following capabilities. Recognizing the intrinsic consistency between pre-training objectives and information compression, researchers began to investigate the relationship between model compression rates and intelligence levels. Wei et al. [56] first reveals the linear correlation between compression ratios and downstream benchmark metrics. Subsequent studies [25] further substantiated the persistence of this linearity across diverse model series and benchmark domains. Despite these advancements, prior work has predominantly treated code as a subdomain of natural language, which overlooks fundamental multi-dimensional distinctions. Code adheres to strict, formal syntactic rules, precluding the ambiguities and flexible grammatical structures inherent in natural language [62]. Furthermore, abstraction in code manifests through explicit, formally-defined constructs such as classes, functions, and APIs, contrasting sharply with the predominantly cognitive and conceptual nature of abstraction in natural language [14, 2]. These critical differences motivate a dedicated inquiry into whether the observed relationship between compression and intelligence in natural language extends uniformly to code.
@@ 2.2 Code Intelligence
The power of LLMs in code-related tasks has spurred the concept of “code intelligence”—the comprehensive ability to understand, generate, and manipulate code [45, 53]. Early evaluations typically focused on Python code completion tasks, such as HumanEval and MBPP [62]. However, as code applications and workflows grow in complexity [55, 49], it is increasingly evident that a robust notion of code intelligence must extend beyond unilingual assessments to encompass proficiency across multiple programming languages [8, 44, 10, 36]. Furthermore, a growing body of research suggests that comprehensive code intelligence is not a monolithic entity but consists of a set of fundamental or “atomic” tasks. While existing taxonomies for code-related LLM tasks vary—for instance, by input/output modalities (e.g., NL-PL, PL-NL) or by user intent [12] —they converge on a core set of underlying capabilities. We classify them to (1) Code Generation: automated creation of source code from natural language descriptions [13, 4, 63]. (2) Code Explanation (or Understanding): grasp and articulate the syntax, semantics, functionality, and underlying intent of a given code snippet [39, 17, 42]. (3) Code Reasoning: infer code properties and predict code behavior without execution [15, 11]. (4) Code Repair: identify and generate patches to correct bugs, vulnerabilities, or other errors in source code [43, 36]. While “code intelligence” is a broad concept, its decomposition into the 4 atomic tasks provides a more structured and actionable framework for quantitative evaluation.
@@ 3 Code Compression as Code Intelligence
We investigate the relationship between compression ratio and code intelligence. We first propose the sliding window BPC for quantifying compression ratio and construct a set of robust validation corpus. Then, we present OmniCodeBench, a multi-lingual and multi-task code benchmark for comprehensive evaluation of code intelligence capabilities of LLMs.
@@ 3.1.1 Compression Ratio Criteria
$BPC=\sum_{i=1}^{N}\log\left(\frac{1}{p_{i}}\right)\cdot\frac{1}{M}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left(-\log(p_{i})\right)}_{\mathrm{train~loss}}\cdot\frac{N}{M}$ (1)
$=R_{1}\cdot R_{2}=R$
where $M$, NN denote corpus size and vocabulary size, respectively. $R_{1}$, the first term of the BPC formula, captures model compression through next-token prediction loss, which is equivalent to the training loss upon first exposure. $R_{2}$, the second term of the BPC formula, represents vocabulary compression rate. It is worth noting that due to differences in vocabulary size, cross-family model comparisons need to consider $R_{1}\cdot R_{2}$, while analysis within the same family can focus solely on $R_{1}$.
@@ 3.1.2 Sliding Window BPC
A critical yet frequently overlooked factor in vanilla BPC evaluation lies in the heterogeneity of models’ context window lengths $W$. Previous practices often simply truncate validation samples exceeding $W$, thereby discarding cross-chunk contextual dependencies and introducing bias when comparing models with divergent context length. To address this, we propose a sliding window protocol with stride adaptation for precise BPC calculation.
@@
Figure 2: Programming language token distribution in the open-source and proprietary validation set. In (b), the unlabeled categories Perl, Shell, C++, Vue, and SQL account for 0.88%, 0.80%, 0.66%, 0.38%, and 0.03%, respectively.
@@ 3.1.3 Robust Validation Corpus Construction
Following the aforementioned criteria, we construct two high-quality corpora, one based on proprietary internal data and the other on publicly available open-source data.
@@
Public corpus   To mitigate linguistic diversity limitation and distributional bias in proprietary datasets, we additionally construct an open-source validation set. We select GitHub as our data source and employ timestamp-based filtering to prevent validation data leakage. Specifically, based on the cutoff date of existing code LLM pre-training data, we collect all GitHub repositories created between May 2024 and November 2024. Subsequently, we follow OpenCoder’s [23] pipeline to perform filtration, removing non-code data and deduplicating the dataset. To minimize bias, we employ rule-based filtering rather than model-based filtering approaches. Additionally, we parse the code to confirm its compilability and apply heuristic-based selection (e.g., filtering by natural language comment length) to preserve high-quality data. However, due to potential historical code leakage from GitHub branches, we adopt MinHash-LSH algorithm to prevent the inclusion of duplicate or near-duplicate code in the validation set. To ensure balanced domain and language distribution in the development set, we perform weighted sampling by repository category, mitigating bias toward high-resource languages. The distribution of validation set is illustrated in Figure 2.
@@ 3.2.1 OmniCodeBench
The evaluation of code intelligence in LLMs remains an open research question. While existing studies predominantly rely on Python-specific benchmarks like HumanEval [13] and MBPP [4], Python alone cannot represent the full spectrum of programming concepts [38]. Our evaluation, therefore, encompasses multiple programming languages. Moreover, conventional assessments focus primarily on the code completion task, which constitutes merely a fraction of real-world coding intelligence. We address this limitation through a comprehensive multi-task evaluation framework. To this end, we propose OmniCodeBench, a multi-lingual and multi-task code benchmark capable of holistic and robustly evaluating code intelligence. Concretely, OmniCodeBench consists of McEval [9] for multi-lingual code generation, BIRD [32] and Spider [59] for SQL code generation, LiveCodeBench [28] for program repair, HumanEvalExplain [41] for code explanation, CRUXEval [20] for code reasoning, and HumanEval [13], MBPP [4], BigCodeBench [63].
@@
StarCoder2-3B 136 1
CodeGemma-2B 55 2
CodeLlama-7B 10 2
Qwen2.5-Coder-7B 13 4
Granite-8B-Code 12 3
@@ 3.2.2 Format Annealing
Code-related tasks require extracting code from the model’s responses and verifying its correctness through test cases. However, the response format of the model is often difficult to perfectly cover, leading to omissions during extraction. Additionally, the model frequently outputs meaningless placeholders, such as todo and pass, severely impacting the accurate evaluation of code intelligence. This phenomenon is caused by the excessive occurrence of related patterns in the pre-training data [23]. To address this, we construct an annealed training set to ensure: (1) preventing external knowledge infusion beyond the original capacity of foundation models, and (2) maintaining strict format compliance with downstream benchmarks. Following these requirements, we integrate three data sources: CommitPackFT [41], CodeSearchNet [27], and LeetCode [23]. All three datasets originate from GitHub, which is the major component of the pretraining corpora of Code LLMs. We run the data cleaning pipelines of various Code LLMs on this data, ensuring the remaining parts were seen during pre-training. Subsequently, all non-functional placeholders are removed, and the content is reformatted according to the prompts of downstream benchmarks, ensuring alignment for evaluation. The final curated dataset contains 4,995 training samples. The complete data synthesis pipeline for each dataset is presented in the supplementary material.
The curated dataset is then used for the specialized continual training phase for all models. During this phase, after each epoch, we leverage the MBPP+ [35] to assess the proportion of post-processed outputs that are empty strings. Training for each model is terminated when this proportion drops to 1%. This early stopping strategy is employed to minimize alterations to the models’ original parameter distributions. Table 1 shows the comparison of the number of empty strings before and after format annealing for five models. It shows that format annealing effectively mitigates the inherent limitations of foundational models in format compliance, ensuring fair comparison through format-controlled evaluation.